# ManiFeel: Benchmarking and Understanding Visuotactile Manipulation Policy Learning

### v1

@@ Abstract
Supervised visuomotor policies have shown strong performance in robotic manipulation but often struggle in tasks with limited visual input, such as operations in confined spaces, dimly lit environments, or scenarios where perceiving the object’s properties and state is critical for task success. In such cases, tactile feedback becomes essential for manipulation. While the rapid progress of supervised visuomotor policies has benefited greatly from high-quality, reproducible simulation benchmarks in visual imitation, the visuotactile domain still lacks a similarly comprehensive and reliable benchmark for large-scale and rigorous evaluation. To address this, we introduce ManiFeel, a reproducible and scalable simulation benchmark for studying supervised visuotactile manipulation policies across a diverse set of tasks and scenarios. ManiFeel presents a comprehensive benchmark suite spanning a diverse set of manipulation tasks, evaluating various policies, input modalities, and tactile representation methods. Through extensive experiments, our analysis reveals key factors that influence supervised visuotactile policy learning, identifies the types of tasks where tactile sensing is most beneficial, and highlights promising directions for future research in visuotactile policy learning. ManiFeel aims to establish a reproducible benchmark for supervised visuotactile policy learning, supporting progress in visuotactile manipulation and perception. To facilitate future research and ensure reproducibility, we will release our codebase, datasets, training logs, and pretrained checkpoints. Please visit the project website for more details: https://zhengtongxu.github.io/manifeel-website/.
Keywords: Tactile sensing, imitation learning, visuotactile policy learning
1 Introduction
@@
In recent years, supervised policy learning has made significant strides in robot manipulation, where visual input is used to generate actions for long-horizon and complex manipulation tasks [1, 2, 3, 4]. However, such vision-based policies encounter notable limitations in environments where visual cues are absent or severely degraded. For example, in cluttered or confined spaces and under low-light conditions, the visual modality may become unreliable, necessitating alternative sensory inputs to ensure robust task execution. In these scenarios, tactile feedback plays a crucial role, compensating for the limitations of vision by providing contact-based sensory information that can effectively guide manipulation.
@@
Combining visual and tactile modalities has emerged as a promising and intuitive approach for learning robotic manipulation, particularly within supervised policy learning frameworks. Recent studies have shown that integrating high-resolution vision-based tactile sensors, such as GelSight [5], can significantly enhance manipulation performance by providing contact-rich information [6, 7, 8, 9, 10]. Various other tactile sensing modalities, such as tactile arrays [11, 12, 13], force-torque sensors [14, 15], tactile 3D point clouds [16], and contact-induced audio signals [17], have also shown strong potential, particularly in tasks requiring fine-grained contact reasoning or operation within unstructured environments.
@@
To address these challenges, we introduce ManiFeel, a scalable and reproducible simulation benchmark for supervised visuotactile policy learning. Our contribution can be summarized as follows.
1) Visuotactile Task Suite: ManiFeel offers a diverse set of tasks and corresponding human demonstration datasets, including insertion, gear assembly, manipulation inside containers, and tasks performed under dim lighting conditions. These tasks are designed to capture a wide range of contact-rich and perception-challenging scenarios, enabling systematic evaluation of how tactile feedback contributes to multimodal policy learning. ManiFeel provides a comprehensive testbed for studying the strengths and limitations of different sensing modalities and promoting the development of better visuotactile policies.
@@
3) Empirical Analysis of Visuotactile Policy Learning: We conduct extensive empirical comparisons on diverse tasks across different policy architectures by systematically varying sensing modalities, tactile representations, and policies. These experiments are conducted in both simulation and real-world settings, allowing us to identify tasks where tactile sensing plays a critical role, uncover effective design choices for visuotactile policy learning, and highlight challenges and promising directions for future research.
@@
4) Real-world Deployment and Reproduction: In addition to constructing simulation benchmarks, we also reproduce the results in real-world settings, comparing different policy architectures and sensing modalities. The consistency between real-world and simulation results further validates the high fidelity of our simulated task suite.
To promote future research, we will open-source the ManiFeel codebase, dataset, training logs, and pretrained checkpoints. We expect ManiFeel to support the visuotactile manipulation community in identifying key challenges, benchmarking progress against prior work, and advancing toward robust solutions for a wide range of visuotactile manipulation tasks across diverse scenarios.
@@
Figure 1: Overview of the manipulation tasks in the ManiFeel benchmark, showing each task’s setup along with the available visual modalities (e.g., wrist camera, agentview camera) and tactile observations (e.g., left and right finger sensors).
2 ManiFeel: Visuotactile Manipulation Learning Benchmark
This section introduces the task suite and modular pipeline design of the ManiFeel benchmark. ManiFeel features a diverse set of manipulation tasks spanning clear-vision conditions to severely degraded or fully occluded visual settings (Section 2.1). The benchmark includes six task categories: tactile sorting in real-world settings, tactile sorting in simulation, tactile exploration and blind insertion in occluded environments, standard insertion tasks, and gear assembly. This results in a total of eight different task setups, including six simulation setups and two real-world setups, as shown in Fig. 1. These tasks are designed to vary in their reliance on visual and tactile modalities, providing a comprehensive framework for analyzing policy performance across different sensory challenges. In addition, we describe the modular pipeline design that enables systematic evaluation of different sensing modalities, tactile representations, and policies across the task suite (Section 2.2).
2.1 Task Suite and Dataset
2.1.1 Task and Environment Design
We design a suite of diverse manipulation tasks to systematically evaluate supervised visuotactile policy learning under varying sensory conditions. The tasks are grouped into six categories, reflecting different levels of reliance on visual and tactile modalities (see Fig. 1). In this section, we provide a high-level overview of each task category. Additional details on task configurations and environment setups are provided in the Appendix B.
Tactile Sorting (Real). We introduce a real-world tactile sorting task involving two spherical objects, a table tennis ball and a golf ball, which are visually similar but have distinct surface textures, making them difficult to distinguish using vision alone. The robot must classify each object and move it into its designated bowl. The task is evaluated under two lighting conditions: normal lighting and dim lighting. In this setting, tactile sensing enables robust object classification based on surface texture, highlighting the complementary role of touch when visual information is degraded or ambiguous.
Tactile Sorting (Simulation). The simulated tactile sorting task mirrors the real-world setup in terms of object types and sensory challenges, but differs in its task objective. Instead of sorting balls into bowls, the robot is required to pick up the designated golf ball and lift it to a specified height. The task is performed under dim lighting conditions to emulate reduced visual quality, emphasizing the importance of tactile feedback for reliable object identification when vision alone is insufficient.
Tactile Exploration (Simulation). In this task, the robot must locate, grasp, and lift a cube-shaped battery, which is randomly placed inside a closed container, to a specified height. The external front-facing camera cannot observe the interior, simulating a vision-occluded environment. The robot relies solely on tactile feedback to explore, localize, and manipulate the object, demonstrating the critical role of touch in environments with severely limited visual information.
Blind Insertion (Simulation). In this task, the robot starts by holding a square peg at an arbitrary orientation and must insert it into a socket located inside a closed container. The external front-facing camera cannot observe the interior, simulating a vision-occluded environment. While the socket pose is fixed, tactile feedback is needed to estimate the in-hand orientation of the peg and to adjust its grasp and alignment for successful insertion under occluded conditions.
Insertion (Simulation). We include insertion tasks that represent fundamental challenges in robotic manipulation, covering standard USB and power plug insertions. The socket location and the in-hand orientation of the objects are randomized at the start of each episode. Successful completion requires leveraging both visual and tactile information to accurately determine the in-hand relative pose and to perform fine adjustments during insertion.
Gear Assembly (Simulation). The gear assembly task simulates an industrial operation where the robot must mount a medium-sized gear onto a central shaft, ensuring correct meshing with side gears. Both the gear base location and in-hand gear orientation are randomized at the start of each episode. The task requires fine spatial alignment, demanding accurate integration of visual and tactile feedback for relative pose estimation and precise adjustment during assembly.
Figure 2: Comparison between tactile images in TacSL simulator [21] and those generated by GelSight Mini. We present examples featuring several objects included in the ManiFeel benchmark.
2.1.2 Observation modalities
ManiFeel is developed on the TacSL simulation platform [21] and supports both vision and tactile observation modalities. For vision, the framework includes multiple camera views and introduces occluded and degraded conditions to emulate realistic manipulation challenges. For tactile sensing, ManiFeel employs a simulated GelSight sensor integrated into a Franka Panda manipulator via TacSL, producing RGB tactile images by measuring surface deformations of a soft elastomer. This setup enables fine-grained perception of object surface geometry during manipulation. Samples of simulated tactile images and their close resemblance to real-world counterparts are shown in Fig. 2, which potentially supports the generalization and applicability of our simulation benchmark to real-world settings. Detailed configurations of camera setups, lighting conditions, and observation modalities for each task are provided in Appendix C.
2.1.3 Teleoperated Demonstration Data Collection
We collect human demonstrations for supervised policy learning using a teleoperation device 3Dconnexion SpaceMouse [^1] https://3dconnexion.com/us/product/spacemouse-compact/buy/. Tactile feedback plays a critical role during data collection, especially for tasks requiring precise contact understanding or operating under visual occlusion and degraded lighting conditions. In such cases, operators rely heavily on real-time tactile observations displayed on a screen to perceive subtle contact events for effective object grasping, classification, and exploration. We argue that access to tactile information during demonstration is crucial for learning effective visuotactile policies. A summary of the collected demonstration data across all tasks is provided in Appendix C. The collected dataset will be open-sourced to support future research on supervised visuotactile policy learning.
2.2 Modular Pipeline Design
In this section, we present the modular pipeline design that forms the basis of our benchmark. The benchmark is designed to systematically evaluate the performance of visuotactile policies which are composed of different sensing modalities, tactile representations, and policies, across a broad set of tasks. This modular structure enables flexible combinations and fair comparisons, ultimately facilitating a deeper understanding of the design principles behind effective visuotactile policy learning. Moreover, the modularity of our design provides strong flexibility for integrating future advancements, such as new tactile modalities (e.g., force-torque sensing), novel tactile representation learning methods, and emerging policy architectures.
Choice of Modalities. To investigate the contribution of different sensing modalities, we evaluate three configurations: vision-only, tactile-only, and visuotactile. This comparison allows us to isolate the role of each modality and assess how they complement each other in different tasks. By varying the available sensory input, we gain insights into when and how tactile sensing improves policy performance, particularly in contact-rich or visually ambiguous scenarios.
Tactile Representations. We benchmark four types of tactile representations: (1) ResNet18 trained from scratch [1, 22], (2) T3 [10], a pretrained encoder using large-scale tactile datasets, (3) UniT [8], a lightweight representation trained with generalizability, and (4) AnyTouch [23], a generalizable tactile representation learned across diverse sensors and modalities. These comparisons aim to reveal whether pretrained or carefully designed tactile encoders offer meaningful advantages in visuotactile policy learning.
Policies. We compare three types of policies: Diffusion Policy [1], Equivariant Diffusion Policy [24], and Flow Matching [25]. These policies differ in their generative modeling approaches and model architectures. Our goal is to understand how each type of policy performs under different input modalities, and whether certain policies are more compatible with the inclusion of tactile sensing. Specifically, we study how tactile signals influence performance across policy architectures.
3 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of diverse visuotactile policies across diverse manipulation tasks using the ManiFeel simulation benchmark and real-world settings. Our study systematically benchmarks different sensory modalities, tactile representations, and policies under varying task conditions within the supervised policy learning framework. Note that all benchmarking modalities, such as vision, tactile, and visuotactile, also include proprioceptive feedback (i.e., the robot end-effector’s pose). For simulation results, success rates are reported as the primary evaluation metric across all experiments. Specifically, success rates are averaged over the last ten epochs of training, across 50 different environment initializations, and over three seeds. This results in each reported success rate being the average of 1,500 evaluations, providing strong statistical significance. For real-world experiments, we perform 15 consecutive rollouts and record the success rate. The initial environment conditions are varied for each rollout. Through extensive experiments, our analysis reveals key factors that influence supervised visuotactile policy learning, identifies the types of tasks where tactile sensing is most beneficial, and highlights key challenges and promising directions for advancing multimodal robotic manipulation.
3.1 Modality Benchmarking
We evaluate the performance of vision-only policies, tactile-only policies, and visuotactile policies across six manipulation tasks using proposed ManiFeel simulation benchmark. All policies use a ResNet18 tactile encoder trained from scratch, and the policy is based on the Diffusion Policy [1], a widely adopted and well-established framework for supervised policy learning. In the experimental setup, external occluded camera views are used for the tactile exploration and blind insertion tasks performed inside a closed container. This setting emulates the limited visual information available to humans when operating in similarly vision-occluded environments and highlights the challenges faced by visual perception in such scenarios. Further details regarding training details and observation configurations are provided in Appendices E–F.
Table 1: Success rates (%) averaged over three seeds across different tasks and sensory modalities.
Modality Tactile Sorting Tactile Exploration Blind Insertion USB Insertion Power Plug Insertion Gear Assembly
Vision-only $30.3\pm 3.0$ $42.9\pm 2.9$ $29.3\pm 1.8$ $51.1\pm 1.7$ $52.3\pm 2.2$ $\mathbf{43.7\pm 3.6}$
Tactile-only $39.0\pm 2.8$ $\mathbf{63.5\pm 2.9}$ $\mathbf{38.7\pm 2.3}$ $2.1\pm 0.8$ $22.5\pm 1.6$ $0.6\pm 0.5$
Visuotactile $\mathbf{60.8\pm 1.9}$ $55.6\pm 3.5$ $31.6\pm 1.6$ $\mathbf{51.4\pm 2.6}$ $\mathbf{55.1\pm 2.8}$ $22.0\pm 2.9$
As shown in Table 1, tactile sensing is critical in tasks where visual information is limited or degraded. Specifically, tactile sorting (conducted under dim lighting), tactile exploration, and blind insertion (both performed inside a closed container) show substantial improvements when using tactile-augmented and visuotactile policies compared to the vision-only baseline. For example, in tactile sorting, success rates improve from 30.3% (vision-only) to 39.0% (tactile-only) and further to 60.8% (visuotactile). Similarly, in tactile exploration, performance increases from 42.9% (vision-only) to 63.5% (tactile-only), although visuotactile performance slightly decreases to 55.6%. In blind insertion, tactile-only achieves 38.7% compared to 29.3% with vision-only, while visuotactile reaches 31.6%. These results highlight the importance of tactile feedback in environments where vision becomes unreliable due to poor illumination or occlusion.
For tactile exploration and blind insertion, the tactile-only policy achieves higher success rates than the visuotactile policy. The inclusion of visual input, heavily occluded by the closed container, appears to introduce uninformative signals that degrade policy learning. This observation highlights a key challenge in multimodal learning: the naive integration of noisy or irrelevant modalities can hinder performance, indicating the need for more selective or adaptive sensory fusion strategies.
For gear insertion and assembly tasks, the performance difference between policies with and without tactile input is less pronounced. In USB insertion, success rates are similar between vision-only and visuotactile policies, which are 51.1% and 51.4%, respectively. Power plug insertion exhibits a similar trend, where tactile sensing provides little additional benefit over vision. In the gear assembly task, however, the visuotactile policy performs worse than the vision-only baseline. These results indicate that vision plays a critical role in these tasks, as visual cues such as object pose and geometry are essential for successful manipulation.
For humans, tactile sensing also plays a crucial role in accomplishing insertion and assembly tasks. However, the results show that simply combining tactile and visual inputs within Diffusion Policy does not guarantee improved performance on these tasks. In gear assembly, in particular, the task complexity arising from the need for precise orientation, teeth alignment, and dynamic interaction with rotating side gears creates a highly challenging multimodal observation space. Naively fusing visual and tactile inputs increases the complexity and variability of the policy input, making direct end-to-end learning more challenging and leading to noticeable performance degradation. These results indicate the need for more effective approaches to visuotactile policy learning that can integrate multimodal information more efficiently and enable more robust policies. Our representation benchmarking (Section 3.2, Table 2) supports this view, showing that UniT [8], a carefully designed pretrained tactile representation, improves gear assembly success by over $12$ percentage points compared to naive fusion. This suggests that learning structured, generalizable visuotactile representations is a promising direction for advancing multimodal policy learning in complex tasks.
Together, these results demonstrate that while tactile sensing is indispensable in visually challenging conditions, its integration must be carefully managed to maximize its benefits and avoid unintended negative impacts on policy performance. Our ManiFeel benchmark provides a structured platform for identifying such key challenges and guiding future research on effective visuotactile policy learning. Additional results, including benchmarking with proprioceptive inputs (e.g., robot proprioception) and comparisons across different camera viewpoints for selective tasks, are provided in Appendix F.
3.2 Tactile Representation Benchmarking
We present simulation benchmarking results for visuotactile policies equipped with different tactile representations, as shown in Table 2. In the experimental setup, visual input is incorporated for the tactile sorting and gear assembly tasks, but omitted for tactile exploration, since including uninformative visual signals caused by the occluded view of the closed container, as discussed in Section 3.1, was found to degrade learning performance. The evaluation is conducted across three representative manipulation tasks: tactile sorting, tactile exploration, and gear assembly. ResNet18 encoder is trained from scratch, while UniT [8], T3 [10], and AnyTouch [23] are pre-trained tactile encoders. Further details on the encoder architectures and training details can be found in Appendices D–E.
Table 2: Success rates (%) averaged over three seeds for each task across various tactile-augmented policy structures of different tactile representations. Bold indicates best performance.
Task ResNet18 UniT T3 AnyTouch
Tactile Sorting $60.8\pm 1.9$ $50.9\pm 2.9$ $52.3\pm 2.1$ $\mathbf{70.4\pm 2.3}$
Tactile Exploration* $\mathbf{63.5\pm 2.9}$ $59.9\pm 1.8$ $59.8\pm 2.8$ $59.8\pm 3.8$
Gear Assembly $22.0\pm 2.9$ $\mathbf{34.5\pm 3.0}$ $25.2\pm 3.0$ $0.9\pm 0.6$
(*) For tactile exploration, the tactile-augmented policy uses only tactile observations without visual input.
AnyTouch achieves the highest success rate on tactile sorting (70.4%) and performs competitively on tactile exploration, demonstrating strong capabilities in tasks that rely on fine contact perception. However, its performance drops considerably on the gear assembly task, suggesting limited generalization to more complex spatial manipulation. In contrast, UniT shows more balanced performance across all tasks, achieving the best result (34.5%) on gear assembly. This highlights the potential of leveraging large-scale pretraining for tactile modalities to enhance supervised visuomotor policy learning. ResNet18, despite being trained from scratch, performs competitively on tactile sorting and tactile exploration tasks, indicating that task-specific training can still be valuable in certain settings. These results suggest that although pre-trained tactile encoders provide strong priors, their effectiveness can vary depending on the specific characteristics of the manipulation task.
Our results further highlight the potential of developing tactile representations that are more specifically tailored to policy learning, aiming for more consistent and superior performance across tasks. Additionally, exploring cross-modal representations that effectively integrate tactile and visual information presents a promising direction for achieving better generalization across diverse domains.
3.3 Policy Benchmarking
In this section, we benchmark three policies: Diffusion Policy (DP) [1], Equivariant Diffusion Policy (EquiDiff) [24], and Flow Matching (FM) [25, 26], across three representative manipulation tasks: tactile sorting, tactile exploration, and gear assembly. Two sensing categories are considered: vision-only policies and tactile-augmented policies where visual input is incorporated for tactile sorting and gear assembly, but not for tactile exploration. For tactile encoder, we use ResNet18 trained from scratch in this experiment. For visual encoder, FM and DP use ResNet18 and EquiDiff uses equivariant ResNet18 [24]. Additional details on the policies and training details can be found in Appendices D–E.
Table 3: Success rates (%) averaged over 3 seeds for each task, comparing vision-only policies and tactile-augmented policies where visual input is incorporated for tactile sorting and gear assembly, but not for tactile exploration*. Results are reported across three different policies. Bold indicates best performance.
Task Vision-only Policy Tactile-augmented Policy
DP EquiDiff FM DP EquiDiff FM
Tactile Sorting $30.3\pm 3.0$ $58.9\pm 3.0$ $5.2\pm 0.6$ $60.8\pm 1.9$ $\mathbf{64.4\pm 3.5}$ $40.9\pm 2.8$
Tactile Exploration* $42.9\pm 2.9$ $43.8\pm 7.8$ $1.2\pm 1.1$ $\mathbf{63.5\pm 2.9}$ $47.1\pm 14.5$ $0.7\pm 0.3$
Gear Assembly $43.7\pm 3.6$ $\mathbf{45.1\pm 3.8}$ $2.5\pm 1.5$ $22.0\pm 2.9$ $42.9\pm 4.8$ $3.3\pm 2.0$
(*) For tactile exploration, the tactile-augmented policy uses only tactile observations without visual input.
As shown in Table 3, tactile-augmented policies consistently outperform vision-only policies across most tasks. Among all policies, EquiDiff achieves the highest overall performance, particularly excelling in tactile sorting and gear assembly tasks. For instance, in the tactile sorting task, EquiDiff improves from 58.9% (vision-only) to 64.4% when tactile input is incorporated, demonstrating its ability to effectively leverage multimodal sensing. In contrast, FM struggles, especially in vision-only policies, suggesting limitations in handling spatial observability challenges. These results demonstrate the effectiveness of the proposed framework for benchmarking and advancing visuotactile policy learning.
3.4 Real-world Experiment
In this section, we evaluate the tactile sorting task in a real-world setting under two lighting conditions: normal and dim. Experiments are conducted using a Franka Panda manipulator [^2] https://franka.de/products/franka-research-3 with a robotic hand integrated with a GelSight Mini [^3] https://www.gelsight.com/gelsightmini/. We benchmark vision-only policies and visuotactile policies with different tactile representations ResNet18 trained from scratch, UniT, and T3. We use Diffusion Policy as the policy. Additional details of the experimental setup can be found in Appendix F.
Table 4: Success rates (%) across 15 trials for the real-world tactile-guided ball sorting task under two lighting scenarios: normal and dim, evaluated across different policy structures. Bold indicates the best performance.
Scenario Vision Visuotactile (ResNet18) Visuotactile (UniT) Visuotactile (T3)
Normal 33.3 86.7 53.3 80.0
Dim 6.7 33.3 40.0 26.7
Table 4 presents the success rates for the real-world tactile-guided ball sorting task under two lighting conditions. Under normal lighting, the visuotactile policy significantly outperforms the vision-only baseline, achieving 86.7% success compared to 33.3%, highlighting the complementary role of tactile feedback even in visually favorable scenarios. The advantage of tactile sensing becomes more pronounced in dimly lit environments, where the performance of vision-only policy drops sharply to 6.7%. In contrast, policies leveraging the tactile modality maintain consistently better performance, with UniT achieving the highest success rate of 40.0%. These results demonstrate the critical importance of tactile sensing for reliable manipulation in visually degraded environments, and validate the effectiveness of learned tactile encoders in enhancing policy robustness. Notably, the relative performance rankings in the normal lighting condition (ResNet18 $>$ T3 $>$ UniT) are consistent between simulation and real-world experiments, suggesting that our proposed simulation benchmark offers a high-fidelity and reliable environment for studying visuotactile manipulation policy learning.
4 Conclusion
This work presents ManiFeel, a simulation benchmark for advancing supervised visuotactile policy learning. Through extensive experiments across diverse manipulation tasks, we systematically evaluate the effects of sensory modalities, tactile representations, and policy structures on performance. Results highlight the critical role of tactile sensing in visually degraded environments and the challenges of integrating tactile inputs in vision-dominant tasks. ManiFeel provides a structured platform for benchmarking visuotactile learning and guiding future research toward effective multimodal integration.
5 Limitation
One limitation of our current benchmark is the lack of support for GelSight marker-based tracking as an input modality for learning contact-rich manipulation policies. Although our benchmark covers a range of visuotactile tasks, it does not yet incorporate observations based on marker displacement, which have recently been shown to play a critical role in enabling tactile-reactive and contact-sensitive policies [9]. Existing simulation methods [21, 27] for marker displacement provide a promising starting point for approximating this modality, and future improvements in the fidelity of such simulations may enable more realistic and informative tactile feedback. Building benchmarks that include simulated marker-based signals and systematically evaluating their impact on policy learning could be a promising direction for future work.
Appendix
Appendix A Related Work
A.1 Imitation Learning
In recent years, imitation learning has progressed rapidly, enabling robots to perform complex, dexterous, and long-horizon manipulation within supervised learning paradigm [1, 2, 3, 4]. Different lines of work have focused on various aspects of imitation learning, such as novel data collection pipelines [3, 4, 28], the design of diverse policy heads such as diffusion [1], flow matching [25], and VQVAE [29], and the development of large-scale policies by integrating vision-language models [26] or scaling up both dataset and model capacity [30, 31].
Traditional imitation learning has predominantly relied on visual modalities, such as images [1, 2, 3] or 3D point clouds [32, 33]. Moreover, visuotactile imitation learning has attracted growing attention. Since manipulation inherently involves interactions with the environment, tactile sensing provides complementary information that is crucial for capturing these interactions. As such, combining visual and tactile modalities to learn manipulation policies via imitation learning has become a natural and promising direction. For instance, several works have leveraged high-resolution vision-based tactile sensors such as GelSight [5], which provide rich contact information and have led to notable improvements in policy learning [6, 7, 8, 9]. In parallel, other tactile modalities, including tactile arrays [11, 12, 13], force-torque sensors [14, 15], tactile 3D point cloud [16], and audio sensors [17], have also demonstrated strong performance, particularly in tasks involving fine-grained manipulation and complex environment interactions.

### v2

@@ Abstract
Supervised visuomotor policies have shown strong performance in robotic manipulation but often struggle in tasks with limited visual inputs, such as operations in confined spaces and dimly lit environments, or tasks requiring precise perception of object properties and environmental interactions. In such cases, tactile feedback becomes essential for manipulation. While the rapid progress of supervised visuomotor policies has benefited greatly from high-quality, reproducible simulation benchmarks in visual imitation, the visuotactile domain still lacks a similarly comprehensive and reliable benchmark for large-scale and rigorous evaluation. To address this, we introduce ManiFeel, a reproducible and scalable simulation benchmark designed to systematically study supervised visuotactile policy learning. ManiFeel offers a diverse suite of contact-rich and visually challenging manipulation tasks, a modular evaluation pipeline spanning sensing modalities, tactile representations, and policy architectures, as well as real-world validation. Through extensive experiments, ManiFeel demonstrates how tactile sensing enhances policy performance across diverse manipulation scenarios, ranging from precise contact-driven operations to visually constrained settings. In addition, the results reveal task-dependent strengths of different tactile modalities and identify key design principles and open challenges for robust visuotactile policy learning. Real-world evaluations further confirm that ManiFeel provides a reliable and meaningful foundation for benchmarking and future visuotactile policy development. To foster reproducibility and future research, we will release our codebase, datasets, training logs, and pretrained checkpoints, aiming to accelerate progress toward generalizable visuotactile policy learning and manipulation.
I Introduction
@@
In recent years, supervised policy learning has made significant strides in robot manipulation, where visual input is used to generate actions for long-horizon and dexterous manipulation tasks [10, 11, 64, 15]. However, vision-based policies face notable limitations not only in environments where visual cues are absent or severely degraded, but also in manipulation scenarios that inherently demand precise contact interactions. In cluttered or confined spaces, under low-light conditions, or during complex contact-rich tasks, relying solely on vision becomes insufficient for robust policy execution. In such contexts, tactile feedback provides essential complementary information, capturing contact geometry, force distribution, and subtle interaction dynamics that vision alone cannot perceive.
@@
Combining visual and tactile modalities has therefore become a promising and intuitive direction for learning robotic manipulation, particularly within supervised policy learning frameworks. Recent studies have shown that integrating high-resolution vision-based tactile sensors, such as GelSight [59], can significantly enhance manipulation performance across diverse contact-rich and visually challenging scenarios [50, 58, 53, 54, 63]. Other tactile sensing modalities, including tactile arrays [17, 30, 4], force-torque sensors [57, 24], tactile 3D point clouds [25], and contact-induced audio signals [31], have likewise demonstrated strong potential for enhancing perception and control in contact-intensive or unstructured environments.
@@
To address these challenges, we introduce ManiFeel, a scalable and reproducible simulation benchmark for supervised visuotactile policy learning. ManiFeel provides a comprehensive testbed for analyzing the strengths and limitations of different sensing modalities, and offering insights into visuotactile policy design. Our contribution can be summarized as follows.
1) Visuotactile Task Suite: ManiFeel offers a diverse set of contact-rich manipulation tasks and corresponding human demonstration datasets, covering insertion, assembly, object exploration inside containers, and sorting tasks under visually challenging conditions. These tasks are designed to capture a wide range of contact-rich and visually limited scenarios, enabling systematic evaluation of how different tactile information contributes to multimodal policy learning.
@@
3) Empirical Analysis of Visuotactile Policy Learning: We conduct extensive empirical comparisons on diverse tasks across different policy architectures by systematically varying sensing modalities, tactile representations, and policies. These experiments are conducted in both simulation and real-world settings, enabling us to identify how different tactile representations support different types of tasks, uncover effective design principles for visuotactile policy learning, and highlight key challenges and promising directions for future research.
@@
4) Real-world Deployment and Reproduction: In addition to constructing simulation benchmarks, we also reproduce the results in real-world settings, comparing different sensing modalities. The correlation between real-world and simulation results further validates the reliability and practical relevance of the ManiFeel simulation benchmark.
5) Open-Source and Reproducible Benchmark Platform: To facilitate reproducible research and accelerate progress in visuotactile manipulation, ManiFeel is fully open-sourced, including its codebase, dataset, and pretrained checkpoints. The benchmark provides a standardized and easily extensible platform that enables fair comparisons and systematic evaluations across different models and sensing configurations. By establishing shared baselines and transparent experimental protocols, ManiFeel lowers the entry barrier for new research and promotes cumulative progress toward more generalizable and robust visuotactile manipulation in diverse real-world scenarios.
@@
Figure 1: Overview of the manipulation tasks in the ManiFeel benchmark, showing each task’s setup along with the available visual modalities (e.g., wrist camera, agentview camera) and tactile observations (e.g., high-resolution tactile images and tactile force-field signals).
II Related Work
II-A Imitation Learning
In recent years, imitation learning has progressed rapidly, enabling robots to perform complex, dexterous, and long-horizon manipulation within supervised learning paradigm [10, 11, 64, 15]. Different lines of work have focused on various aspects of imitation learning, such as novel data collection pipelines [64, 15, 48], the design of diverse policy heads such as diffusion [10], flow matching [61, 62], and VQVAE [29], and the development of large-scale policies by integrating vision-language models [6] or scaling up both dataset and model capacity [46, 65].
Traditional imitation learning has predominantly relied on visual modalities, such as images [10, 11, 64] or 3D point clouds [60, 55, 26, 70]. Moreover, visuotactile imitation learning has attracted growing attention. Since manipulation inherently involves interactions with the environment, tactile sensing provides complementary information that is crucial for capturing these interactions. As such, combining visual and tactile modalities to learn manipulation policies via imitation learning has become a natural and promising direction. For instance, several works have leveraged high-resolution vision-based tactile sensors such as GelSight [59], which provide rich contact information and have led to notable improvements in policy learning [50, 58, 53, 54]. In parallel, other tactile modalities, including tactile arrays [17, 30, 4], force-torque sensors [57, 24], tactile 3D point cloud [25], and audio sensors [31], have also demonstrated strong performance, particularly in tasks involving fine-grained manipulation and complex environment interactions.